### v2

# JLQCD collaboration

Abstract
We investigate spatial two-point correlation functions of mesonic operators in two-flavor lattice QCD at high temperatures. The simulated temperatures cover the range $T\in[147,330]$ MeV, where the critical temperature is estimated around 165 MeV. To ensure a good control of the chiral symmetry we employ the Möbius domain-wall fermion action for two degenerate flavors of quarks. With a lattice cut off $a^{-1}\sim 2.6$ GeV, the residual mass is reduced to 0.14 MeV. With the energy spectrum obtained from the screening mass at incremental values of the temperature range, we examine the $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry, the anomalous axial $U(1)$ as well as an enhanced symmetry which exchanges the spin degrees of freedom. We also study how the data approaches the perturbative prediction given by twice the Matsubara frequency of free quarks.
preprint: OU-HET-1250
I Introduction
The chiral phase transition of quantum chromodynamics (QCD), which occurred in the early universe, gives a crucial input for understanding how hadrons formed and acquired their $O(1)$ GeV masses we observe currently. While it is naturally assumed that the lightest two quarks flavors play a crucial role in the phase transition, the details of the critical phenomena depend on the symmetry of the up and down quarks above the critical temperature $T_{c}$ Pisarski:1983ms .
In the massless limit of the up and down quarks, it is natural to assume that the standard $SU(2)_{L}\times SU(2)_{R}$ symmetry is recovered at some critical temperature in the range 150–200 MeV and no other symmetry exists. However, the axial $U(1)_{A}$, which is broken by anomaly, may effectively appear if nontrivial topological excitations of gluons are strongly suppressed above $T_{c}$. When such topological effects are described by instantons Gross:1980br ; Diakonov:1984vw ; Morris:1984zi ; Schafer:1996wv ; Ringwald:1999ze , which is valid for pure Yang-Mills, QCD with heavy quarks, and QCD with a large number of color degrees of freedom, strong suppression of the quantum anomaly is not expected and the axial $U(1)_{A}$ is likely to remain broken at $T\sim T_{c}$. However, with the light dynamical quarks, the typical size of the topological excitation is described by the pion physics whose correlation length becomes larger than $1/T$,and this description of the topological excitations by instantons is no longer valid. The fate of the axial $U(1)_{A}$ anomaly at high $T$ has been discussed in both theoretical works Cohen:1996ng ; Cohen:1997hz ; Aoki:2012yj ; Pelissetto:2013hqa ; Kanazawa:2014cua ; Sato:2014axa ; Nakayama:2016jhq as well as numerical works Cossu:2013uua ; Buchoff:2013nra ; Dick:2015twa ; Brandt:2016daq ; Ishikawa:2017nwl ; Bazavov:2019www .
Another interesting possibility is that at high $T$, QCD may in fact have a larger symmetry. This symmetry is related to the anti-periodic boundary condition of quarks imposed in the temporal direction. At very high temperature, the mesons would have large masses proportional to the temperature $T$, due to the Matsubara mass. As a consequence, a nontrivial symmetry may appear, analogous to the heavy quark symmetry Neubert:1993mb ; Brambilla:2004jw at $T=0$. In fact, in the perturbative evaluation in Laine:2003bd the masses of all the mesons of different spins converge to $2\pi T$. Note that this is not a symmetry of the original QCD Lagrangian but is an “emergent” symmetry in the large $T$ limit. Although this is only an approximate symmetry within the perturbative framework in the large $T$ limit, nonperturbative gluonic effects may enhance this symmetry even at $T\sim 2T_{c}$, this symmetry is called the chiral-spin symmetry Glozman:2014mka ; Glozman:2016swy ; Glozman:2017dfd ; Lang:2018vuu ; Catillo:2021rrq ; Glozman:2022lda .
It is, therefore, a crucial task to investigate the symmetry at high temperatures through a nonperturbative approach to QCD, which must include a consistency check with the perturbative prediction of the screening masses converging to $2\pi T$. Lattice QCD simulations that have previously studied this issue, however, used fermion formulations that explicitly broke the $SU(2)_{L}\times SU(2)_{R}$ symmetry. This has made the analysis ambiguous in identifying the symmetry breaking/restoration patterns apart from lattice artifacts. In particular, we reported in Cossu:2015kfa ; Tomiya:2016jwr ; Aoki:2020noz that the lattice artifacts in the probes of axial $U(1)_{A}$ anomaly are enhanced at high temperatures and it is difficult to disentangle the physical effect from the discretization errors.
In this work, we employ the domain-wall fermion formalism Kaplan:1992bt ; Shamir:1993zy ; Furman:1994ky with an improvement proposed in Refs.Brower:2005qw ; Brower:2012vk . With this Möbius domain-wall fermion, which is a good approximation of the overlap fermion Neuberger:1997fp , we achieve a theoretically clean control of the symmetry of QCD. With a fixed lattice spacing $a\sim 0.075$ fm, the residual mass, which is an indicator of the explicit symmetry violation is reduced to 0.14(6) MeV. In addition, we set the quark masses so that the lightest bare quark mass $ma=0.001$ is less than the physical mass value of the up and down quarks. The simulated temperatures span a range $T\in[147,330]$ MeV, which corresponds to $[0.9,2]T_{c}$, where the $T_{c}$ is estimated from the peak of the chiral susceptibility Aoki:2024uvl . The main runs are carried out with a fixed lattice size $L=32$ except for the lowest two temperatures, where the pions behave as the pseudo-Nambu-Goldstone bosons; in this case we simulate two additional volumes, $L=40$ and $48$, to control the finite volume systematics.
On each ensemble of O(1000-10000) trajectories, we measure the two-point mesonic correlation functions in 6 different iso-triplet channels every 50 trajectories. With the screening masses obtained from fitting the correlators, we examine the symmetry at each simulated temperature and quark mass. In our main analysis of the mesonic correlators, the standard $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry is examined by the difference between the vector $V$ and axial vector $A$ correlators. The tensor $T_{t}$ and $X_{t}$ channels (definitions will be given below), rather than scalar $S$(noisy) and pseudoscalar $PS$ channels, are useful to estimate how much the axial $U(1)_{A}$ anomaly effect remains.
Because of the Matsubara mass, the system may exhibit the chiral-spin $SU(2)_{CS}$ symmetry mentioned above. In our previous work at higher temperatures Rohrhofer:2017grg ; Rohrhofer:2019qwq ; Rohrhofer:2019qal we observed the emergence of this $SU(2)_{CS}$ symmetry from the correlator ratios among $V$, $A$, $T_{t}$ and $X_{t}$. A similar observation simulating quarks with a good chiral symmetry was reported in Refs. Chiu:2023hnm ; Chiu:2024jyz ; Chiu:2024bqx . In this work, we revisit this enlarged symmetry group, computing the screening masses and simulating lower temperatures down to $T\sim 147$ MeV, which is $\sim 0.9T_{c}$.
The rest of the paper is organized as follows. In Sec. II, we disucuss the general functional form of the mesonic correlators and the possible pattern of the symmetry at high temperatures. In particular, we address the enhancement of symmetry compared to the original QCD using the effective theory of heavy quarks due to the Matsubara mass. Our lattice setups as well as our methodology for how to extract the screening mass are addressed in Sec. III. Numerical results for our lattices are given in Sec. IV. And finally, we give a summary and discuss our numerical results in Sec. V. Some results in this work were already given in Ward:2024wze and preliminary results for $N_{f}=2+1$ simulations were presented in Ward:2024tdm . We also refer the readers to Refs. Aoki:2021qws ; JLQCD:2024xey ; Goswami:2024kcq for the related works from the JLQCD finite $T$ project.
II Mesonic two-point functions and symmetries at high temperatures
In the limit of massless up and down quarks, the QCD action is invariant under the standard chiral transformation of the group $SU(2)_{L}\times SU(2)_{R}$, as well as that of the axial $U(1)_{A}$. The axial $U(1)_{A}$ symmetry is in general broken by quantum anomaly, but its magnitude at finite temperatures is nontrivial since it is tightly related to the topological excitations of gluons, which may be strongly suppressed at high temperatures.
In this work, we consider the mesonic two-point function in the $z$ direction whose continuum limit is given by
[…]
where the quark bilinear operator $O_{\Gamma}=\bar{q}\Gamma T^{a}q$ is taken to be a flavor-triplet (the isospin generator is denoted by $T^{a}$). For $\Gamma$ we choose $\gamma_{5}(PS),\mathbb{I}(S),\gamma_{1,2}(V),\gamma_{5}\gamma_{1,2}(A),\gamma_{4}\gamma_{3}(T_{t})$ and $\gamma_{5}\gamma_{4}\gamma_{3}(X_{t})$. The list of operators we measure and the symmetries connecting them are summarized in Tab. 1.
Under the standard chiral $SU(2)_{L}\times SU(2)_{R}$ rotation:
$q\to\exp(i\alpha_{a}T^{a}+i\beta_{a}T^{a}\gamma_{5})q,\;\;\;\bar{q}\to\bar{q}\exp(-i\alpha_{a}T^{a}+i\beta_{a}T^{a}\gamma_{5}),$ (2)
only $V$ and $A$ correlators mix while all the other channels remain unchanged up to isospin exchanges. Therefore, the $SU(2)_{L}\times SU(2)_{R}$ symmetry breaking can be detected by the difference between the $V$ and $A$ correlators.
For the axial $U(1)_{A}$ transformation,
$q\to\exp(i\epsilon\gamma_{5})q,\;\;\;\bar{q}\to\bar{q}\exp(i\epsilon\gamma_{5}),$ (3)
[…]
$\Gamma$ Reference Name Abbr. Symmetry Correspondences
$\mathbb{I}$ Scalar S }$\bigg{\}}U(1)_{A}$
$\gamma_{5}$ Psuedo Scalar PS
$\gamma_{1},\gamma_{2}$ Vector $V$$\bigg{\}}SU(2)_{L}\times SU(2)_{R}$
[…]
$\gamma_{4}\gamma_{3}$ Tensor $T_{t}$$\bigg{\}}U(1)_{A}$
$\gamma_{4}\gamma_{3}\gamma_{5}$ Axial Tensor $X_{t}$
Table 1: List of quark bilinear operators we compute the two-point correlation functions and the symmetries connecting them.
In addition to the above manifest symmetries in the QCD Lagrangian, we may expect emergence of larger symmetries due to anti-periodic boundary condition on the quarks in the temporal direction. This enhancement of symmetry is analogus to the heavy quark symmetry Neubert:1993mb ; Brambilla:2004jw , which (approximately) appears in the effective theory of bottom or charm quarks.
In order to see this enhanced symmetry, let us consider a free quark propagator in the positive $z$ direction:
$\langle q(z)\bar{q}(0)\rangle(p_{1},p_{2})$ $=T\sum_{p_{0}}\int\frac{dp_{3}}{(2\pi)}\frac{e^{ip_{3}z}}{i\gamma_{0}p_{0}+i\gamma_{3}p_{3}+i\gamma_{1}p_{1}+i\gamma_{2}p_{2}+m}$
$=T\sum_{p_{0}}\int\frac{dp_{3}}{(2\pi)}\frac{-(i\gamma_{0}p_{0}+i\gamma_{3}p_{3}+i\gamma_{1}p_{1}+i\gamma_{2}p_{2}-m)e^{ip_{3}z}}{p_{0}^{2}+p_{3}^{2}+p_{2}^{2}+p_{1}^{2}+m^{2}}$
$=T\sum_{p_{0}}\frac{-(i\gamma_{0}p_{0}-\gamma_{3}E(p_{0},p_{1},p_{2})+i\gamma_{1}p_{1}+i\gamma_{2}p_{2}-m)e^{-E(p_{0},p_{1},p_{2})z}}{2E(p_{0},p_{1},p_{2})}$ (4)
where $E(p_{0},p_{1},p_{2})=\sqrt{p_{0}^{2}+p_{1}^{2}+p_{2}^{2}+m^{2}}$ and the other components $p_{1},p_{2}$ are fixed. Here the Matsubara mass $p_{0}=(2n+1)\pi T$ takes discrete values due to the antiperiodicity in the imaginary temporal direction. In the large $T\gg m,p_{1},p_{2}$ limit, the lowest $p_{0}=\pm\pi T$ will dominate the signal and the correlator becomes
$\langle q(z)\bar{q}(0)\rangle(p_{1},p_{2})$ $=T\left[\gamma_{3}\frac{1+i\text{sgn}(p_{0})\gamma_{0}\gamma_{3}}{2}e^{-\pi Tz}+O(1/T)\right].$ (5)
In the same way, the correlator in the temporal direction with a fixed $p_{1},p_{2},p_{3}$ in the large Matsubara mass limit is expressed as
$\langle q(t)\bar{q}(0)\rangle(p_{1},p_{2},p_{3})$ $=-iT\left[\sum_{p_{0}}\frac{\gamma_{0}p_{0}}{p_{0}^{2}}\exp(ip_{0}t)+O(1/T)\right].$ (6)
These propagators are invariant under the following transformation:
$q$ $\to\exp(i\Sigma^{i}\epsilon_{i})q,$
$\bar{q}$ $\to\bar{q}\gamma_{0}\exp(-i\Sigma^{i}\epsilon_{i})\gamma_{0},$ (7)
where $\Sigma^{i}$ has three components,
$\Sigma^{i}=(\gamma_{k},-i\gamma_{5}\gamma_{k},\gamma_{5}),\;\;\;k=1,2.$ (8)
We do not distinguish $k=1$ and $k=2$, since they are identical under the rotational symmetry along the $z$ axis. The generators $\Sigma^{i}$ form an $SU(2)$ algebra, which is identical to the chiral-spin $SU(2)_{CS}$ symmetry suggested in the literature Glozman:2014mka ; Glozman:2016swy ; Glozman:2017dfd ; Lang:2018vuu ; Catillo:2021rrq ; Glozman:2022lda .
Note that any correction from gauge field potential only appears in the next-to-leading $1/T$ contribution. Thus, the chiral spin $SU(2)_{CS}$ symmetry emerges at sufficiently high temperatures. With this $SU(2)_{CS}$ symmetry, the mesonic operators $(A_{1},T_{t},X_{t})$ form a nontrivial triplet. These operators have different spins in the original four dimensions, but they are all in the spin-one representation of $SO(3)$ symmetry along the $z$ axis. One can consider another set $(V_{2},PS,S)$ but the operators transform differently under the $SO(3)$ rotation and therefore, the mass degeneracy would not be that good111 We thank L. Glozman for pointing out the difference of the two triplets..
From (5), it is obvious that in the free quark limit the meson screening mass is $2\pi T$. In fact, in the perturbative analysis Laine:2003bd the screening mass is obtained as
$M_{\Gamma}=2\pi T+cg^{2}T,$
with a numerical constant $c$ coming from the one-loop correction where $g$ is the gauge coupling. It is remarkable that the result does not depend on $\Gamma$. This result is analogous to the heavy quark symmetry Neubert:1993mb ; Brambilla:2004jw which is an approximate symmetry of the bottom or charm quarks at $T=0$. As the spin and orbital angular momentum dependent interactions are proportional to the inverse of the heavy quark mass, the spectrum of the quarkonia and heavy-light mesons are insensitive to spin. The same may explain here that with a large Matsubara mass proportional to $T$, the operators $(V_{1},PS,S)$ with different spin form an $SU(2)_{CS}$ triplet222 The $SO(3)$ rotational symmetry is broken also among $(A_{1},T_{t},X_{t})$ at very high temperatures, since the temporal direction shrinks as $1/T\to 0$. However, its effect is sub-leading. . In Ref. Chiu:2023hnm , it was suggested that other combinations of tensors $\gamma_{1}\gamma_{3}(T_{k})$ and $\gamma_{1}\gamma_{3}\gamma_{5}(X_{k})$ form a quartet with $A_{4}$ and $V_{4}$. Inclusion of this quartet into our analysis is of interest to future work; however, in this paper we will focus on the triplet.
It is an important numerical subject to quantify the emergent $SU(2)_{CS}$ symmetry at intermediate temperatures $T\sim T_{c}$ and see if it has any impact on the chiral phase transition. It is also interesting to investigate if an additional transition exists at $T>T_{c}$, triggered by this emergent symmetry. These three symmetries, accessible through the differences between mesonic screening masses from the pairs of the triplet channels333 In our previous study Rohrhofer:2019qal , we also studied the ratio of differences of correlators $|C_{A}(z)-C_{T}(z)|/|C_{A}(z)-C_{S}(z)|$. However, we find that the denominator approaches to zero at high temperatures (presumably due to the large Matsubara mass and the analogy to “heavy quark symmetry” we discussed in Sec. II.) and the ratio is not a good indicator of the $SU(2)_{CS}$ symmetry. In this work, we focus on the screening mass differences and present the values in the physical unit (MeV) so that one can easily compare with the simulated temperature for each set of data. , constitute our motivation to simulate QCD at high temperatures in the range $[0.9,2]T_{c}$, for a range of quark masses covering the physical point as a method to understand the symmetry structure of QCD at high temperatures.
III Lattice setups
For gauge configuration generation, we perform hybrid Monte Carlo simulations [^4] See Refs. Cossu:2013ola ; Boyle:2015tjk ; Ueda:2014rya ; Amagasa:2015zwb for details of our simulation codes. with the tree-level improved Symanzik gauge action Luscher:1985zq and the Möbius domain-wall fermion action Brower:2005qw ; Brower:2012vk . See Ref. Tomiya:2016jwr for details of our simulations. In this work, we focus on the gauge configurations at a fixed lattice spacing $0.075$ fm generated with the bare coupling constant $\beta=4.30$. A portion of the same gauge link ensembles are shared with our previous studies Aoki:2021qws ; Aoki:2020noz , in addition to these we also use an updated and newly generated series of ensembles on larger lattices at lower temperatures. The simulation parameters are summarized in Tab. 2.
For the scale setting, we perform the Wilson flow and use the reference flow time $t_{0}=(0.1539\;\mbox{fm})^{2}$ presented in Sommer:2014mea to determine the lattice cut-off $2.6431/a=$ 2.643 GeV. Our main lattice size $L=32$ corresponds to 2.37 fm, while we also simulate $L=40$ and 48 lattices for the lowest two simulated temperatures to study the finite size systematics near the critical temperature.
We vary the temperature by changing the temporal size of the lattice from 8 to 18, which corresponds the range 147 MeV $\leq T\leq$ 330 MeV. The critical temperature at the physical point is estimated to be $T_{c}\sim 165(3)$ MeV from the peak position of the disconnected chiral susceptibility Aoki:2024uvl where the error is statistical only.
The lightest bare sea quark mass $m=0.001$ corresponds to 2.6 MeV, which is estimated to be 71% of the physical quark mass $m_{\text{phys}}=0.0014(2)$.
For the link variables in the Dirac operator, we perform the stout smearing Morningstar:2003gk three times with the smoothing parameter $\rho=0.1$. The size of the extra dimension for the Möbius domain-wall fermion action is set to $L_{s}=16$. With this treatment, the residual mass or the chiral symmetry breaking of the lattice Dirac operator is suppressed to 0.14(6) MeV.
For the correlator measurement, we use the four-dimensional effective Dirac operator,
$D_{\text{DW}}^{\text{4D}}(m)=\frac{1+m}{2}+\frac{1-m}{2}\gamma_{5}\tanh(L_{s}\tanh^{-1}(H_{M})),$ (9)
where the Möbius kernel $H_{M}$ is
$H_{M}=\gamma_{5}\frac{2D_{W}}{2+D_{W}},$ (10)
and $D_{W}$ is the standard Wilson-Dirac operator with a negative mass $M=-1$. Here and in the following, we set $a=1$ unless explicitly specified.
For each ensemble, 40-700 measurements were performed per 50 trajectories and our numerical data, after binning between 2-10 samples, did not show any significant autocorrelations. Using the rotational symmetry of the z-directional correlators, we applied averaging over three different spatial directions. For our sources, we did not apply any additional source smearing and used point-like source and sink operators.
IV Numerical results
Figures 13 present the typical effective mass plots with the $\cosh$ fitting ansatz at the lightest quark mass $m=0.001$ at the lowest three simulated temperatures $T=$ 147, 165 and 190 MeV  [^5] We also present the corresponding raw correlators in the Appendix.. It is not difficult to find a reasonable plateau, in contrast to the study with staggered fermions in Ref. Bazavov:2019www , where a sizable contamination from the excited states was reported. A possible reason for such a difference is the fact that the Möbius domain-wall fermion is free from the unphysical taste degrees of freedom as well as unwanted operator mixing among different multiplets of the $SU(2)_{L}\times SU(2)_{R}$ flavor group. From the data in the fitting range, which is shown by the shadow in these plots, we determine the meson screening masses for the different channels. The obtained values and their differences among various symmetry multiplets are summerized in Tab. 2 and 3.
In Fig. 4, we plot the temperature dependence of the obtained meson screening masses at our lightest mass $m=0.001$ ($\sim$ 2.6 MeV). As disscused in Sec. II, the data at high $T$ for all channels approach $2\pi T$, which is twice of the lowest mode Matsubara mass, shown by the grey line. At lower temperatures around or below the critical temperature $T_{c}\sim 165$(3) MeV, shown by the vertical grey band, there is a drastic change in the flucutation of the masses. The scalar mass, in particular, shows inflated errorbars at the lower temperatures, while quickly converging to a less noisy result, close to the $PS$ channel data at higher temperatures. For reference, we put the experimental results of $\pi^{\pm}$, $\rho$, $a_{0}$ and $a_{1}$ mesons at zero temperature, indicated by colored bars corresponding to the associated channels on the vertical axis. It is interesting to note that the lowest temperature, $T=147$ MeV data are already quite close to the experimental values at $T=0$; this may indicate that the chiral symmetry breaking by the quark condensate is already strong enough to form hadrons. In the $T=0$ limit, the $T_{t}$ and $X_{t}$ two-point functions transform in the same way as $V$ and $A$ correlators, respectively, and share the same intermediate states.
The same plot but with the temperature normalized by $T_{c}=165$ MeV and the screening mass normalized by $2\pi T$ is shown in Fig. 5. For a comparison, the results by HotQCD collaboration Bazavov:2019www which simulated $N_{f}=2+1$ HISQ quarks are shown by the shadows(normalized by $2\pi T$ and the 2+1-Flavor $T_{c}$ estimate respectively). Although the value of $T_{c}$ is different due to the absence/presence of the strange quark, the two results appear consistent and share many of the same features. The only exception is the $S$ data at the lowest two simulated temperatures; as explained in Bazavov:2019www , the shadow approaches the twice of the pion mass in the $T=0$ limit, due to a lattice artifact of the staggered quark action. In our Möbius domain-wall fermion formulation, this never happens since the transition from the $S$ triplet to two $PS$ mesons is prohibited by the exact isospin symmetry.
In Fig. 7, we plot the mass difference between $A$ and $V$ channels $\Delta m_{V-A}$ , which probes the $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry breaking, as a function of temperature $T$. Here the data from the largest spatial volume for each simulated temperature is plotted. Around the critical temperature shown by the vertical grey band, the mass difference $\Delta m_{V-A}$ quickly vanishes. This is most clearly seen at our lightest mass $m=0.001$, we find $\Delta m_{V-A}/m_{A}\sim 0.1$% at $T=190$ MeV and 2.8% at $T=165$ MeV666 In recent study Gavai:2024mcj with Möbius domain-wall quarks, violation of the axial $U(1)_{A}$ symmetry even at $T=186$ MeV was reported. We notice, however, that their lattice spacings $a>0.13$ fm are coarser than the one $a\sim 0.1$ fm we simulated in our previous work Cossu:2015kfa where we find a larger violation of the Ginsparg-Wilson in the low-lying eigenmodes of the Dirac operator than the residual mass, which leads to overestimation of the axial $U(1)_{A}$ anomaly. . Both values are consistent with zero to within a standard deviation.
A similar behavior but with larger errorbars is found in Figs. 7 and 9 where the mass differences between $X_{t}$ and $T_{t}$, $\Delta m_{X-T}$, and that of $PS$ and $S$, $\Delta m_{PS-S}$, are plotted. These are the probes for the axial $U(1)_{A}$ symmetry. We find that $\Delta m_{X-T}/m_{X}$ is 2.5% at $T=190$ MeV and 12.7% at $T=165$ MeV. The latter is consistent with zero within a larger errorbar. This indicates that the chiral anomaly effect is quite suppressed above $T_{c}$.
Fig. 9 is a plot of the chiral-spin $SU(2)_{CS}$ symmetry, which is an approximate symmetry expected up to $O(1/T)$ corrections, as explained in Sec. II. This symmetry is examined by the mass difference $\Delta m_{X-A}$ between $X_{t}$ and $A$. The $T$ dependence for the screening mass, in addition to the reduction in noise, is analogous to $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$; however, we find a qualitative difference of $SU(2)_{CS}$ in Fig. 10. Here we plot the quark mass dependence of $\Delta m_{V-A}$ (squares), $\Delta m_{PS-S}$ (circles), $\Delta m_{X-T}$ (upward triangles), and $\Delta m_{X-A}$ (downward triangles) at $T=330$ MeV. The former three probing $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries, are consistent with zero, and their uncertainty is $\sim 1$MeV, which is less than 1% of $T$. In contrast to these other mass differences, the data $\Delta m_{X-A}$ is clearly nonzero even at the lightest mass: $\sim-40$ MeV showing that the associated chiral-spin symmetry $SU(2)_{CS}$ is only approximate: $|\Delta m_{X-A}|/T\sim 0.12$. We note that at much higher temperatures than $T=1$ GeV, a better degeneracy even between $PS$ and $V$ channels was reported in Ref. DallaBrida:2021ddx .
Finally let us discuss possible systematics in our numerical results. In Fig. 11, we plot the lattice size $L$ dependence of all measured screening masses at $T=147$, 165 and 220 MeV. The data at different $L$ are consistent within two standard deviations and we do not see any significant finite volume dependence. Due to our fixed lattice spacing at $a=0.075$ fm, we cannot numerically estimate the discretization effects. Our choice is relatively fine compared to other groups and we expect the automatic $O(a)$ improvement due to the chiral symmetry of the Möbius domain-wall fermion action. Note also, that our simulated up and down quark mass range covers the physical point, and the lightest quark mass is sufficiently close to the chiral limit.
Figure 1: Effective screening mass plots at $T=147$ MeV. The data for $PS,S,A,T_{t}$ channels at the lightest quark mass $m=0.001$ are presented.
Figure 2: The same plots as Fig. 1 but at $T=165$ MeV.
Figure 3: The same plots as Fig. 1 but at $T=190$ MeV.
Figure 4: $T$ dependence of the screening masses for various meson channels at the lightest simulated mass $m=0.001$($\sim$2.6 MeV). The grey line shows $2\pi T$, which is twice of the Matsubara mass. For a reference, we put the experimental results of $\pi^{\pm}$, $\rho$, $a_{0}$ and $a_{1}$ mesons at zero temperature, indicated by bars at the vertical axis. The vertical grey band indicates our estimate for the critical temperature $T_{c}=165(3)$ MeV from the chiral susceptibility. Note that the busy data points are slightly shifted horizontally.
Figure 5: The same figure as Fig. 4 but the temperature is normalized by $T_{c}=165$ MeV. and the screening mass is normalized by $2\pi T$. For a comparison, the results by HotQCD collaboration Bazavov:2019www which simulated $N_{f}=2+1$ HISQ quarks are shown by the shadows. Except for the $S$ channels at the lowest two simulated temperatures, the two results look consistent with each other.
Figure 6: Temperature $T$ vs. the screening mass difference between $V$ and $A$ which probes the $SU(2)_{L}\times SU(2)_{R}$ symmetry. The vertical grey band indicates our estimate for the critical temperature $T_{c}=165(3)$ MeV from the chiral susceptibility.
Figure 7: Temperature $T$ vs. the screening mass difference between $X_{t}$ and $T_{t}$ which probes the $U(1)_{A}$ symmetry. The vertical grey band indicates our estimate for the critical temperature $T_{c}=165(3)$ MeV from the chiral susceptibility.
Figure 8: Temperature $T$ vs. the screening mass difference between $PS$ and $S$ which probes the $U(1)_{A}$ symmetry. The vertical grey band indicates our estimate for the critical temperature $T_{c}=165(3)$ MeV from the chiral susceptibility.
Figure 9: Temperature $T$ vs. the screening mass difference between $A_{x}$ and $X_{t}$ which probes the $SU(2)_{CS}$ symmetry. The vertical grey band indicates our estimate for the critical temperature $T_{c}=165(3)$ MeV from the chiral susceptibility.
Figure 10: Quark mass dependence of $\Delta m_{V-A}$ (squares), $\Delta m_{PS-S}$ (circles), $\Delta m_{X-T}$ (upward triangles), and $\Delta m_{X-A}$ at $T=330$ MeV. Compared to the former three probing $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries, which are consistent with zero with errorrbars $\sim 1$ MeV, the chiral-spin probe $\Delta m_{X-A}$ (downward triangles) is clearly nonzero: $\sim-40$ MeV and the associated chiral-spin symmetry $SU(2)_{CS}$ is only approximate: $|\Delta m_{X-A}|/T\sim 0.12$.
Figure 11: The lattice size $L$ vs. screening masses at $T=147$, 165 and 220 MeV is plotted. We do not see any significant finite volume dependence beyond $2\sigma$.
V Summary and discussion
We have simulated $N_{f}=2$ lattice QCD at finite temperatures in the range $[0.9,2]T_{c}$ with the Möbius domain-wall quark action and Symanzik gauge action setting the lattice spacing at $a\sim 0.075$ fm. Our simulated quark mass covers the physical point, and thus the chiral symmetry is well under control with the residual mass $\sim 0.14$ MeV. For the lowest two temperatures, we have simulated multiple volumes in order to check the finite size systematics.
The two-point correlation functions of the isovector operators with different spins have been measured in the spatial directions. From the standard $\cosh$ fit, we have extracted the screening masses for the six channels of the isospin triplet. The temperature dependence of the screening masses shows several remarkable features. First, the values at the lightest simulated quark mass $m=0.001$ at $T=0.9T_{c}$ are consistent with the experimental values at $T=0$. This may indicate that the chiral symmetry is strongly broken, even for high temperatures below the psuedocritical point. Second, the scalar $S$, axial vector $A$ and axial tensor $X_{t}$ masses all reduce above $T\sim T_{c}$ to values similar to their $SU(2)_{L}\times SU(2)_{R}$ or $U(1)_{A}$ partners $PS$, $V$ and $T_{t}$. Then above $T_{c}$, all the channels monotonically increase and appear to converge to twice the ground state Matsubara mass $2\pi T$.
Taking the difference of the screening masses between the aforementioned channel partners, we have examined various symmetries of QCD at high temperatures. The standard chiral $SU(2)_{L}\times SU(2)_{R}$ is restored at the critical temperature $T_{c}\sim 165$MeV (observable from the degeneracy of the $V$ and $A$ channels), consistent with the critical temperature estimated from the chiral susceptibility in previous work. The axial $U(1)_{A}$ symmetry behaves similarly, with degeneracy between the $T_{t}$ and $X_{t}$ as well as $PS$ and $S$ at or close to $T_{c}$ although the signals are noisier. The mass difference between these three channel pairs at $T\sim 2T_{c}$ is consistent with zero with quite small uncertainty less than 1 MeV, or 1% of the temperature. In contrast to the above two symmetries, the chiral spin $SU(2)_{CS}$ probed by $A$ and $X_{t}$ is only approximate at $T\sim 2T_{c}$, showing a breaking $\sim 40$ MeV, which is 12% of the temperature.
Acknowledgements.
We thank T. Cohen, H.-T. Ding, C. Gattringer, L. Glozman, A. Hasenfratz, C.B. Lang, O. Phillipsen, R. Pisarski, S. Prelovsek, for fruitful discussions. Discussions during the long-term workshop, HHIQCD2024, at Yukawa Institute for Theoretical Physics (YITP-T-24-02) were useful to complete this work. We also thank the members of JLQCD collaboration for their encouragement and support. Numerical simulations are performed on IBM System Blue Gene Solution at KEK under a support of its Large Scale Simulation Program (No. 16/17-14), Oakforest-PACS at JCAHPC under a support of the HPCI System Research Projects (Project IDs: hp170061, hp180061, hp190090, hp200086, and hp210104) and Multidisciplinary Cooperative Research Program in CCS, University of Tsukuba (xg17i032 and xg18i023), the supercomputer Fugaku provided by the RIKEN Center for Computational Science (hp200130, hp210165, hp210231, hp220279 and hp230323), Wisteria/BDEC-01 Odyssey at JCAHPC (HPCI: hp220093, hp230070 and hp240050 MCRP: wo22i038), and Polarie and Grand Chariot at Hokkaido University (hp200130). We used Japan Lattice Data Grid (JLDG) for storing a part of the numerical data generated for this work. This work is supported in part by the Japanese Grant-in-Aid for Scientific Research (No. JP26247043, JP18H01216 and JP18H04484, JP22H01219), and by MEXT as “Priority Issue on Post-K computer” (Elucidation of the Fundamental Laws and Evolution of the Universe) and by Joint Institute for Computational Fundamental Science (JICFuS).
Figure 12: Plots of the raw correlators for $T=147$MeV and $165$MeV. All plot of the correlators correlators are averaged over the three spatial directions and symmetrized with respect to the reflection of the axis $z\to-z$.
Figure 13: The three higher temperature raw correlator plots like fig.12. These plots are also symmetrized around the midpoint of the correlator.
[…]
$48^{3}\times 18$ 3.6 147 2.7 0.00100 2.6 146 113(34) 951(111) 1106(182) 1120(201) 1068(154)
0.00250 6.6 40 1097(288) 220(5) 1009(69) 1404(129) 1801(166) 1012(71)
0.00375 9.9 40 1387(337) 245(5) 854(74) 1145(155) 1279(223) 1164(97)
0.00500 13.2 83 344(86) 135(7) 903(58) 1375(177) 1733(281) 1199(117)
$36^{3}\times 18$ 2.7 147 2.0 0.00100 2.6 146 166(22) 812(85) 941(132) 1483(206) 1391(115)
0.00250 6.6 121 2605(533) 220(12) 767(53) 930(156) 1645(153) 1228(117)
0.00375 9.9 122 1586(333) 240(11) 886(77) 1046(181) 1635(107) 1216(52)
0.00500 13.2 131 1490(139) 288(6) 861(43) 1231(125) 1527(117) 1091(56)
$40^{3}\times 16$ 3.0 165 2.5 0.00100 2.6 165 642(251) 149(46) 1109(165) 1141(185) 994(268) 1453(305)
0.00250 6.6 95 750(412) 206(17) 544(146) 536(201) 1499(196) 1161(146)
0.00375 9.9 97 1831(145) 301(16) 1003(90) 1400(158) 1156(305) 1364(171)
0.00500 13.2 95 3093(336) 273(20) 936(113) 1330(216) 1007(341) 721(95)
$32^{3}\times 16$ 2.4 165 2.0 0.00100 2.6 165 233(49) 883(93) 890(84) 1485(256) 1633(222)
0.00250 6.6 116 2252(272) 276(25) 904(75) 978(133) 1268(431) 1212(176)
0.00375 9.9 163 1624(538) 309(14) 934(90) 1047(131) 1552(178) 966(152)
0.00500 13.2 143 444(187) 326(24) 932(84) 1106(159) 1679(207) 1216(89)
$32^{3}\times 14$ 2.4 189 2.3 0.00100 2.6 190 381(40) 353(39) 998(40) 998(41) 1187(62) 1161(61)
0.00250 6.6 177 1608(287) 401(27) 1150(49) 1164(52) 1376(154) 1407(131)
0.00375 9.9 137 368(76) 410(25) 964(33) 996(44) 1318(146) 1212(135)
0.00500 13.2 133 1262(261) 409(29) 940(44) 1069(74) 1385(114) 1178(82)
$48^{3}\times 12$ 3.6 220 4.0 0.00100 2.6 220 793(43) 792(43) 1306(37) 1306(37) 1377(39) 1378(39)
0.00250 6.6 97 913(21) 911(20) 1290(39) 1293(39) 1411(38) 1410(38)
0.00375 9.9 114 1033(192) 706(47) 1189(78) 1235(93) 1441(48) 1431(54)
0.00500 13.2 116 888(108) 752(109) 1234(33) 1236(35) 1378(58) 1316(48)
$40^{3}\times 12$ 3.0 220 3.3 0.00500 13.2 220 1534(351) 502(51) 1327(32) 1298(28) 1291(116) 1425(38)
0.01000 26.4 244 1495(336) 696(35) 1222(24) 1275(32) 1302(106) 1444(99)
$32^{3}\times 12$ 2.4 220 2.7 0.00100 2.6 532 712(61) 701(58) 1281(38) 1283(39) 1402(58) 1408(58)
0.00250 6.6 534 697(88) 797(85) 1235(22) 1234(23) 1294(46) 1307(51)
0.00375 9.9 689 615(158) 696(77) 1244(23) 1251(24) 1570(103) 1374(104)
0.00500 13.2 544 1350(360) 717(90) 1182(25) 1213(23) 1465(115) 1292(69)
0.01000 26.4 622 1254(226) 686(36) 1227(57) 1160(76) 1345(101) 1302(80)
$24^{3}\times 12$ 1.8 220 2.0 0.00100 2.6 373 747(84) 741(82) 1283(27) 1280(29) 1427(28) 1434(28)
0.00250 6.6 361 1637(334) 806(121) 1292(30) 1295(30) 1345(77) 1305(80)
0.00375 9.9 331 687(120) 785(53) 1333(32) 1335(33) 1392(32) 1398(32)
0.00500 13.2 363 2505(488) 736(134) 1219(31) 1245(36) 1327(66) 1469(55)
0.01000 26.4 365 770(153) 713(54) 1271(29) 1307(38) 1421(33) 1450(29)
$32^{3}\times 10$ 2.4 264 3.2 0.00500 13.2 640 1319(35) 1266(26) 1566(18) 1566(18) 1666(32) 1631(26)
0.00800 21.1 237 1252(19) 1251(19) 1572(15) 1574(15) 1626(19) 1627(20)
0.01000 26.4 291 1525(121) 1117(94) 1560(19) 1581(27) 1630(49) 1654(28)
0.01500 39.6 121 1458(123) 1236(100) 1596(21) 1599(21) 1666(28) 1693(29)
$32^{3}\times 8$ 2.4 330 4.0 0.00100 2.6 260 1810(11) 1809(11) 2026(9) 2026(9) 2083(13) 2083(13)
0.00500 13.2 317 1796(19) 1791(18) 2038(12) 2038(12) 2094(24) 2094(24)
0.01000 26.4 350 1783(15) 1781(15) 2033(8) 2033(8) 2082(9) 2082(9)
0.01500 39.6 306 1828(24) 1791(22) 2027(10) 2028(10) 2064(17) 2063(17)
0.02000 52.9 218 1807(15) 1796(14) 2013(18) 2014(18) 2069(15) 2066(15)
0.04000 105.7 164 1824(27) 1770(20) 2021(16) 2024(16) 2080(16) 2073(15)
Table 2: Simulation parameters and meson screening mass in the physical unit [MeV] obtained from the $\cosh$ fit.
[…]
$48^{3}\times 18$ 3.6 147 2.7 0.00100 2.6 146 -155(120) 52(233) -8(196)
0.00250 6.6 40 -877(289) -395(119) 789(201) -398(158)
0.00375 9.9 40 -1141(335) -291(155) 115(274) -133(201)
0.00500 13.2 83 -209(88) -472(179) 534(381) -361(297)
$36^{3}\times 18$ 2.7 147 2.0 0.00100 2.6 146 -129(66) 92(290) -539(213)
0.00250 6.6 121 -2385(525) -163(139) 417(235) -710(169)
0.00375 9.9 122 -1346(331) -160(171) 419(127) -586(114)
0.00500 13.2 131 -1203(141) -370(127) 437(147) -296(144)
$40^{3}\times 16$ 3.0 165 2.5 0.00100 2.6 165 -493(272) -32(75) -459(536) 148(377)
0.00250 6.6 95 -544(417) 8(121) 338(307) -963(184)
0.00375 9.9 97 -1530(142) -397(143) -208(447) 244(329)
0.00500 13.2 95 -2820(337) -394(279) 287(304) 323(439)
$32^{3}\times 16$ 2.4 165 2.0 0.00100 2.6 165 -7(28) -149(464) -595(281)
0.00250 6.6 116 -1975(279) -74(151) 56(542) -290(445)
0.00375 9.9 163 -1315(539) -113(128) 586(247) -505(248)
0.00500 13.2 143 -118(172) -174(138) 463(221) -573(227)
$32^{3}\times 14$ 2.4 189 2.3 0.00100 2.6 190 -28(14) 0(4) 26(16) -188(67)
0.00250 6.6 177 -1208(294) -14(16) -30(274) -213(159)
0.00375 9.9 137 42(55) -32(35) 107(265) -321(140)
0.00500 13.2 133 -853(240) -129(75) 207(144) -315(127)
$48^{3}\times 12$ 3.6 220 4.0 0.00100 2.6 220 -2(1) -0(0) -1(0) -71(37)
0.00250 6.6 97 -2(1) -3(1) 0(1) -118(47)
0.00375 9.9 114 -327(206) -46(35) 10(57) -206(79)
0.00500 13.2 116 -136(196) -2(13) 62(86) -141(62)
$40^{3}\times 12$ 3.0 220 3.3 0.00500 13.2 220 -1032(399) 14(18) -22(96) -141(98)
0.01000 26.4 244 -799(347) -48(20) -2(79) -87(79)
$32^{3}\times 12$ 2.4 220 2.7 0.00100 2.6 532 -11(7) -2(2) -5(4) -119(80)
0.00250 6.6 534 100(102) 1(2) -13(32) -60(36)
0.00375 9.9 689 81(109) -6(2) 197(195) -320(98)
0.00500 13.2 544 -633(444) -31(17) 173(135) -252(112)
0.01000 26.4 622 -568(243) 67(79) 44(131) -185(124)
$24^{3}\times 12$ 1.8 220 2.0 0.00100 2.6 373 -6(4) 4(4) -7(5) -148(38)
0.00250 6.6 361 -831(225) -3(2) 40(67) -48(97)
0.00375 9.9 331 98(74) -1(1) -6(6) -56(43)
0.00500 13.2 363 -1769(375) -26(15) -144(111) -81(78)
0.01000 26.4 365 -57(114) -36(23) -28(36) -115(45)
$32^{3}\times 10$ 2.4 264 3.2 0.00500 13.2 640 -53(53) -1(1) 35(18) -100(28)
0.00800 21.1 237 -0(3) -1(1) -0(3) -53(20)
0.01000 26.4 291 -407(203) -21(26) -25(67) -49(69)
0.01500 39.6 121 -222(217) -3(1) -27(28) -67(30)
$32^{3}\times 8$ 2.4 330 4.0 0.00100 2.6 260 -1(1) -0(0) -0(0) -57(11)
0.00500 13.2 317 -5(5) -0(0) 0(0) -56(26)
0.01000 26.4 350 -2(1) -0(0) 0(0) -49(10)
0.01500 39.6 306 -37(35) -1(0) 1(1) -37(13)
0.02000 52.9 218 -12(6) -2(0) 2(1) -54(12)
0.04000 105.7 164 -54(30) -4(2) 6(3) -55(11)
Table 3: The difference between the screening masses of the mesons in the physical unit [MeV].
Appendix A Raw correlators
In this appendix, we summarize the results for the raw correlator with the lightest quark mass $m=0.001(2.6\mbox{MeV})$ at the simulated temperatures. Some data having negative values are neglected.